when displacement is one half of amplitude. ratio kinetic to potential energy?
 

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,

If the question is of SHM which is highly likely to be, then the answer is
Potential energy=U=kx2/2 (x is the displacement)
Total energy= T=kA2/2 (A is the amplitude)
U=kA2/8
T=kA2/2
Kinetic energy=K=3kA2/8

Ratio=K/U
=3:1
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)